### v3

# Each generic polytope in $\mathbb{R}^{3}$ has a point with ten normals to the boundary

Abstract.
It is conjectured since long that each smooth convex body $\mathbf{P}\subset\mathbb{R}^{n}$ has a point in its interior which belongs to at least $2n$ normals from different points on the boundary of $\mathbf{P}$. The conjecture is proven for $n=2,3,4$.
We treat the same problem for convex polytopes in $\mathbb{R}^{3}$ and prove that each generic polytope has a point in its interior with $10$ normals to the boundary.
This bound is exact: there exists a tetrahedron with no more than $10$ normals emanating from a point in its interior.
Key words and phrases:
Morse theory, bifurcation, critical point, binormal.    MSC 52B70
1. Introduction
It is conjectured since long that for any smooth convex body $\mathbf{P}\subset\mathbb{R}^{n}$ there exists a point in its interior which belongs to at least $2n$ normals from different points on the boundary of $\mathbf{P}$. The conjecture is a simple exercise for $n=2$. There are two proofs for $n=3$: a geometrical proof by E. Heil [3], and a topological proof by J. Pardon [6].
So, each $3$-dimensional smooth body has a point with at least $6$ emanating normals. This is an exact bound: an ellipsoid (with all different lengths of axes) has at most $6$ normals from a point in its interior.
For $n=4$, the conjecture was proven by J. Pardon [6], and (to the best of our knowledge) nothing is known for higher dimensions.
In this paper we solve a similar problem for compact convex polytopes with non empty interior.
Definition 1.
Let $\mathbf{P}\subset\mathbb{R}^{3}$ be a compact convex polytope with non empty interior, let $y\in Int\leavevmode\nobreak\ \mathbf{P}$. Let a support plane $h$ to $\mathbf{P}$ and a point $z\in h\cap\mathbf{P}$ be such that the line $yz$ is orthogonal to $h$. Then $yz$ is called a normal to the boundary of $\mathbf{P}$ emanating from $y$. The point $z$ is called the base of the normal.
Our main result is:
Theorem 1.
Each generic [^1] Genericity is defined in Definition 2. polytope in $\mathbb{R}^{3}$ has a point with (at least) $10$ emanating normals to the boundary.
Three necessary remarks are:
  • In the smooth case the bases of the normals to $\partial\mathbf{P}$ emanating from $y$ are the critical points of the squared distance function
    $SQD_{y}(x)=|x-y|^{2}:\partial\mathbf{P}\rightarrow\mathbb{R}.$
    Generically, $SQD_{y}$ is a Morse function, thus Morse theory gives a tool. With some efforts, Morse-theoretic machinery extends also to polytopes [5]. So counting normals to the boundary is equivalent to counting critical points of $SQD_{y}$.
  • Each polytope in $\mathbb{R}^{3}$ has a point with at least $8$ normals to the boundary. This follows straightforwardly from Morse counts, see [5]: let $y$ be the center of the sphere of maximal radius contained in $\mathbf{P}$. Then the function $SQD_{y}$ has at least four local minima (the tangent points) and at least one maximum, therefore at least three saddles.
  • The bound in Theorem 1 is exact: there exists a tetrahedron with no more than $10$ normals from a point in its interior [5].
2. Definitions and preliminaries
Most of the definitions and lemmata of the section are borrowed from [5]; we repeat them here for the sake of completeness.
2.1. Active regions, bifurcation set and its sheets
Let $\mathbf{P}\subset\mathbb{R}^{3}$ be a compact convex polytope with non-empty interior.
Let $F$ be a face, an edge, or a vertex of a polytope $\mathbf{P}$. The active region $\mathcal{AR}(F)$ is the set of all points $y\in Int\,\mathbf{P}$ such that $F$ contains the base of some normal emanating from $y$.
The bifurcation set $\mathcal{B}(\mathbf{P})$ is the union of the boundaries of active regions. It is the PL counterpart of the focal set that appears in the smooth case.
The bifurcation set is piecewise linear. Its parts not lying on $\partial\mathbf{P}$ (sheets) fall into two types, red and blue, see Fig. 1.
Figure 1. Active regions of a face and an edge
That is,
  1. (1)
    The active region of a face $F$ is the intersection of the infinite right angled prism based on $F$ and the polytope $\mathbf{P}$.
    The side faces of the prism are called blue sheets of $\mathcal{B}(\mathbf{P})$.
  2. (2)
    The active region of an edge $E$ sharing two faces $F_{1}$ and $F_{2}$ is the intersection of the polytope $\mathbf{P}$ with the wedge-like polyhedron bounded by two planes containing $E$ and orthogonal to $F_{1}$ and $F_{2}$ (these are blue sheets), and another two planes containing the endpoints of $E$ and orthogonal to $E$ (red sheets).
  3. (3)
    The active region of a vertex $V$ is the inner normal cone of $\mathbf{P}$ at the vertex $V$ (bounded by red sheets) intersected with the polytope $\mathbf{P}$.
Although the function $SQD_{y}$ is non-smooth, it can be treated as a Morse function. So counting critical points of $SQD_{y}$ amounts to counting normals, exactly as it is in the smooth case.
Lemma 1.
For $y\notin\mathcal{B}(\mathbf{P})$, we have:
  1. (1)
    Local maxima of $SQD_{y}$ are attained at some vertices of $\mathbf{P}$.
  2. (2)
    Local minima are attained at some faces of $\mathbf{P}$.
  3. (3)
    Saddles are attained at some edges of $\mathbf{P}$.
Lemma 2.
If $y$ crosses transversally one sheet of $\mathcal{B}(\mathbf{P})$, a pair of critical points of $SQD_{y}$ either dies or is born.
  1. (1)
    Crossing a blue sheet amounts to birth-death of a minimum and a saddle.
  2. (2)
    Crossing a red sheet amounts to birth-death of a maximum and a saddle.
We say that a point $y\in Int(\mathbf{P})$ projects to a face $F$ (to an edge $e$, respectively) if the orthogonal projection of $y$ to the affine hull of $F$ (of $e$, respectively) lies in the interior of $F$ (of $e$).
An edge of a polytope with an acute dihedral angle is called an acute edge.
Lemma 3.
  1. (1)
    A point $y\in\mathbf{P}$ projects to the interior of a face iff the face contributes a minimum of $SQD_{y}$.
  2. (2)
    Let $e$ be an acute edge of $\mathbf{P}$. Then a point $y\in\mathbf{P}$ projects to $e$ iff the function $SQD_{y}$ attains a saddle at the edge $e$.
Lemma 4.
For a convex polytope $\mathbf{P}\subset\mathbb{R}^{3}$ and a point $y\notin\mathcal{B}(\mathbf{P})$, the number of critical points of $SQD_{y}$ (equivalently, the number of normals emanating from $y$) equals $2+2s$, where $s$ is the number of saddles.
In particular, the number of normals emanating from $y$ is even.
In the paper we restrict ourselves by generic polytopes:
Definition 2.
A polytope $\mathbf{P}$ with the vertex set $\mathbf{V}$ is generic if affine hulls of any two (two-element or three-element) subsets of $\mathbf{V}$ are neither parallel nor orthogonal, unless this is dictated by unavoidable reasons. For example,
i:
No two edges are orthogonal.
ii:
No edge is parallel to a face, unless they are incident.
Important observations are: (1) For a generic polytope, the sheets of $\mathcal{B}(\mathbf{P})$ intersect transversally. (2) Any polytope can be turned to a generic one by a close to $id$ projective transform. (3) A sheet of $\mathcal{B}$ related to a face contains no vertices of $\mathbf{P}$ except for vertices of the face.
2.2. Associated spherical geometry
A spherical polygon is a polygon with geodesic edges lying in the standard sphere $S^{2}$. It is always supposed to fit in an open hemisphere.
A vertex $V$ of a polytope $\mathbf{P}$ yields a convex spherical polygon $Spher(V)$ which is the intersection of a small sphere centered at $V$ with the polytope $\mathbf{P}$. Although the sphere is small, let us assume that it is equipped with the metric of the unit sphere $S^{2}$.
For points $X,Y\in S^{2}$ denote by $|XY|$ the spherical distance between them.
Say that a point $Y$ projects to a geodesic segment $s$ if (i) the segment and the point $Y$ fit in an open hemisphere, (ii) $s$ contains a point $W$ such that the geodesic segment $YW$ is orthogonal to $s$, and (iii) $|YW|<\pi/2$. In this case $|YW|$ realizes the Hausdorff distance between the point $Y$ and the segment $s$.
Let $P\subset S^{2}$ be a convex spherical polygon, let $Y\in Int\leavevmode\nobreak\ P$. Define the spherical squared distance function
$SQD_{Y}^{sph}:\partial P\rightarrow\mathbb{R},\ \ SQD_{Y}^{sph}(Y)=|XY|^{2}.$
Let $V$ be a vertex of $\mathbf{P}$. Let a point $y$ lie in the interior of $\mathbf{P}$, close to the vertex $V$. Here and in the sequel we denote by $Y$ the intersection of the small sphere centered at $V$ and the ray $Vy$.
Definition 3.
The core $Core(P)$of a convex spherical polygon $P$ is the set of its interior points $Y$ such that the distance from $Y$ to each point from $\partial P$ does not exceed $\pi/2$.
Lemma 5.
In the above notation,
  1. (1)
    $V$ is a local maximum of $SQD_{y}$ iff $Y\in Core(Spher(V))$.
  2. (2)
    Let $e$ be an edge of $\mathbf{P}$ incident to $V$. Assume $A=Spher(V)\cap e$. The edge $e$ contributes a saddle point of $SQD_{y}$ iff
    1. (a)
      $A$ is a local maximum of $SQD_{Y}^{sph}$, and
    2. (b)
      $|AY|<\pi/2$.
  3. (3)
    A face $F$ incident to $V$ contributes a minimum of $SQD_{y}$ iff $Y$ projects to the edge $F\cap S^{2}$ of $Spher(V)$.
Definition 4.
  1. (1)
    A point $Y\in Int(P)$ has a short maximum (respectively, short minimum) at $X\in\partial P$ if $X$ is a local maximum (respectively, a local minimum) of $SQD^{sph}_{Y}$, and $|YX|<\pi/2$.
  2. (2)
    A convex spherical polygon $P\subset S^{2}$ is nice if there is a point $Y$ in its interior such that $SQD^{sph}_{Y}$ has at least three short maxima. Otherwise $P$ is called skew.
  3. (3)
    A vertex $V$ of polytope $\mathbf{P}\subset\mathbb{R}^{3}$ is skew if the associated spherical polygon $Spher(V)$ is skew.
Short maxima are attained at (some) vertices, whereas short minima are attained at edges. (This is no longer true for non-short maxima and minima.)
For a fixed $Y$, maxima and minima of $SQD_{Y}^{sph}$ alternate on $\partial P$. We say that a minimum and a maximum are neighbours if there is no other maxima and minima on a connecting arch of $\partial P$.
Example. Skew triangles were studied in [5] in details. A skew triangle is depicted in Fig 2.
Figure 2. A skew triangle
3. Proof of Theorem 1
Proposition 1.
If a generic polytope $\mathbf{P}\subset\mathbb{R}^{3}$ has no point in its interior with 10 emanating normals, then all the vertices of $\mathbf{P}$ are skew.
Proof.
Assume the contrary: $V$ is a nice vertex of $\mathbf{P}$, and $\mathbf{P}$ has no point in its interior with 10 emanating normals. Let $Y$ be a point inside $P=Spher(V)$ with three short maxima $A_{1},A_{2},A_{3}$. Each of the maxima is surrounded on $\partial P$ by two short minima, so there are at least three short minima of $SQD^{sph}_{Y}$. Consider two cases.
  1. (1)
    The number of the short minima (that are neighbors of $A_{i}$) is exactly $3$. Then all the minima are short, $Y\in Core(Spher(V))$, so $Y\in\mathcal{AR}(V)$.
    Therefore, for every point on the ray $VY$, the vertex $V$ is a local maximum. Take a point $y$ lying on the ray very close to $V$ and let it travel along the ray. We may assume that $y$ never crosses two sheets of the bifurcation set at a time. If this is not the case, perturb $y$ a little bit.
    Initially $SQD_{y}$ has three minima $m_{1}$, $m_{2}$, and $m_{3}$, lying on the faces incident to $V$ , three saddles $s_{1}$, $s_{2},s_{3}$, lying on edges that correspond to $A_{1},A_{2},A_{3}$, and one local maximum $V$. Since $V$ is not the global maximum, there is at least one other maximum. Altogether there are at least 8 critical points of $SQD_{y}$, or, equivalently, 8 normals emanating from $y$.
    The point $y$ keeps moving until the first event, which is either a bifurcation of $SQD_{y}$, or the point $y$ leaves the interior of $\mathbf{P}$. We have the following case analysis:
    (a) The first event is a birth of two new critical points. This gives us 10 normals.
    (b) The point $y$ leaves the polytope $\mathbf{P}$. This means that $y$ hits one of the faces which is not incident to the vertex $V$. Right before $y$ leaves the polytope $\mathbf{P}$, there is the fourth minimum on this face, and we have at least 10 critical points.
    (c) One of the saddles $s_{i}$ dies. This happens in two cases: (i) $s_{i}$ meets a (local) maximum of $SQD_{y}$. The maximum cannot be $V$ since $s_{i}$ slides away from $V$. It is not any other maximum $V^{\prime}$. Indeed, recall that $s_{i}$ is the orthogonal projection of $y$ to the edge. If $s_{i}$ meets $V^{\prime}$ before $y$ leaves $P$, then $|Vy|>|VV^{\prime}|$. Then $V^{\prime}$ can’t be maximum of $SQD_{y}$ when $y=V$. (ii) $s_{i}$ meets a minimum. It can be none of $m_{1}$, $m_{2}$, and $m_{3}$, so right before the bifurcation, the function $SQD_{y}$ has $4$ minima and $2$ maxima, which proves the claim.
  2. (2)
    The number of short minima that are neighbors of $A_{i}$ is at least $4$.
    We repeat the above construction. Initially we have at least four minima of $SQD_{y}$, three saddles, and one maximum. The bifurcations are analyzed in the same way. If there is no bifurcation until $y$ leaves $\mathbf{P}$, we arrive at a fifth minimum just before leaving the polytope, which gives $10$ normals.
The following properties of skew polygons are crucial for the proof:
Proposition 2.
  1. (1)
    Each skew polygon has exactly two acute angles, say, at vertices $M_{1}$ and $M_{2}$,
  2. (2)
    $|M_{1}M_{2}|>\pi/2$,
  3. (3)
    For any $p\in P$, the inequalities $|M_{1}p|<\pi/2$ and $|M_{1}p|<\pi/2$ imply that $p\in Core(P)$.
We postpone the proof of the proposition to the next section.
Proof of Theorem 1 modulo Proposition 2
By Proposition 2, (1) each vertex of $\mathbf{P}$ is incident to exactly two acute edges. Therefore, acute edges form a closed broken line on the boundary of $\mathbf{P}$. Denote the broken line by $\mathcal{A}$.
For an interior point $y\in\mathbf{P}$ set
$SQD_{y,\mathcal{A}}:\mathcal{A}\rightarrow\mathbb{R},\ \ SQD_{y,\mathcal{A}}(x)=|x-y|^{2}.$
Lemma 6.
  1. (1)
    Minima of $SQD_{y,\mathcal{A}}$ are attained at edges of $\mathcal{A}$. Each such a minimum is a saddle critical point of
    $SQD_{y}:\partial P\rightarrow\mathbb{R}.$
  2. (2)
    Maxima of $SQD_{y,\mathcal{A}}$ are attained at vertices of $\mathcal{A}$. Maxima of $SQD_{y,\mathcal{A}}$ are exactly maxima of $SQD_{y}$.
Proof.
(1) An edge contains a minimum of $SQD_{y,\mathcal{A}}$ iff $y$ projects to it. Since the edge is acute, by Lemma 6 we have a saddle of $SQD_{y}$.
(2) One way is clear: a maximum of $SQD_{y}$ is also a maximum of $SQD_{y,\mathcal{A}}$. The converse follows from Proposition 2, (3) and Lemma 5.
Definition 5.
A binormal of $\mathbf{P}$ is a segment $VW$ with endpoints on $\partial\mathbf{P}$ such that there exist two parallel support planes orthogonal to $VW$, one passing through $V$, and the other one through $W$.
Lemma 7.
There exists a binormal $VW$ of $P$ such that $V$ is a vertex, and $W$ lies in the interior of some edge.
Proof.
Bimormals of $\mathbf{P}$ correspond to critical points of the width function of $\mathbf{P}$.  [^2] The width function is clearly not a smooth function. So our aguments rely on the concept of Clarke subdifferential borrowed from non-smooth analysis, see e.g. [1], which contains also the definition of critical point and the regular interval theorem.
We have the following case analysis:
(1) If the endpoints of a binormal are two vertices of $\mathbf{P}$, then the binormal realises a local maximum of the width function.
(2) If the endpoints of a binormal are a vertex of $\mathbf{P}$ and a point in the interior of a face, then the binormal realises a local minimum.
(3) If the endpoints of a binormal lie in the interiors of some (non-parallel, because of genericity) edges of $K$, then the binormal also realises a local minimum.
Since the width function is an even function defined on $S^{2}$, or equivalently, a function defined on the projective plane, it cannot have maxima and minima only. Therefore there should be other types of binormals. The only remaining generic type is the one we need. ∎
Now we are ready to prove the theorem.
Consider a binormal from Lemma 7. Let its endpoints be $V$ (a vertex of $\mathbf{P}$) and $W$ (a point in the interior of an edge). Remind that by assumption and Proposition 1, $V$ is a skew vertex (as well as all the other vertices of $\mathbf{P}$.
Case 1. The binormal $VW$ lies in the interior of $\mathbf{P}$, Fig. 3.
Observe that $VW$ belongs to the active region $\mathcal{AR}(V)$, that is, for every point on the binormal, $V$ is a local maximum.
We may assume that the binormal never crosses two different sheets of $\mathcal{B}(\mathbf{P})$ at a time. If this is not the case, replace the binormal by a generic broken line emanating from $V$ and ending at $W$ lying close to the binormal.
Take a point $y$ lying on the binormal very close to the vertex $V$. By Lemmata 5 and 3, $SQD_{y}$ has two saddles $s_{1},s_{2}$ on the acute edges incident to $V$, a saddle $s_{3}=W$, two minima $m_{1},m_{2}$ on the faces incident to $V$, and a local maximum at $V$. Since $V$ is not a global maximum, there is some other global maximum; altogether there are at least $8$ normals emanating from $y$.
Figure 3. Binormal $VW$
Lemma 8.
Let $VV^{\prime}$ and $VV^{\prime\prime}$ be two acute edges incident to $V$. For a point $y$ lying close to $V$ none of $V^{\prime}$ and $V^{\prime\prime}$ is a (local) maximum of $SQD_{y}$.
Proof.
By Proposition 2, (2), there is an edge incident to $V^{\prime}$ which makes an obtuse angle with $VV^{\prime}$. Therefore $V^{\prime}$ is not a maximum neither for $V$ nor for any point $y$ close to $V$. ∎
Now let the point $y$ travel along the binormal starting at $V$. As we have seen, initially there are at least emanating $8$ normals.
Keep moving until the first event, which is either a bifurcation (that is, a crossing of a sheet of $\mathcal{B}(P)$), or the point $y$ reaches $W$. We have the following case analysis.
  1. (1)
    The first event is a birth of a new pair of critical points. Then right after bifurcation we have at least 10 normals.
  2. (2)
    The first event is a death of two critical points.
    1. (a)
      One of the saddles $s_{i}$ meets a local maximum, and the two critical points die. This saddle cannot be $s_{3}$ since $s_{3}$ does not move. It can be neither $s_{1}$ nor $s_{2}$ by Lemma 8.
    2. (b)
      One of the saddles $s_{1,2}$ meets a minimum. The minimum point is necessarily attained at a face incident to $V$, which is impossible since the distances between the bases of the normals grow as $y$ slides away from $V$.
    3. (c)
      The saddle $s_{3}$ meets a minimum. This happens only when $y$ reaches $W$. This case is analyzed below.
  3. (3)
    Point $y$ reaches $s_{3}=W$ without bifurcation. Once $y$ almost reaches $W$, we observe two new minima lying on faces incident to the edge containing $W$. By condition, no face of $\mathbf{P}$ contains both $V$ and $W$, therefore none of the two new minima coincides with $m_{1}$ or $m_{2}$. Since there were no bifurcations on the way, all the ’’old’’ critical points persist, and eventually we have $4$ minima, $2$ maxima, and $3$ saddles. Since the number of critical points is even, altogether there are at least 10 normals.
Case 2. $VW$ lies in the interior of a face of $\mathbf{P}$.
In this case the edge containing $W$ is acute.
One repeats the arguments from Case 1 for a point $y$ travelling along the binormal shifted a little bit inwards the polytope. The same bifurcation arguments survive: if $y$ crosses a sheet of $\mathcal{B}(P)$ on its way, we get $10$ emanating normals.
So the only remaining case is the one when there are no bifurcations on the (shifted) binormal. Here we need additional arguments since one of the two new minima coincides with either $m_{1}$ or $m_{2}$. However we conclude that for a point $y$ on the (shifted) binormal we have (at least) $3$ minima, $2$ maxima, and $3$ saddles. Moreover, these $3$ saddles are attained on acute edges of $P$. That is, we have three minima of $SQD_{y,\mathcal{A}}$, see Lemma 6. Since $\mathcal{A}$ is a topological circle (or a disjoint union of topological circles), there are necessarily three maxima of $SQD_{y,\mathcal{A}}$. By Lemma 6, they are maxima of $SQD_{y}$ as well.
Now for $y$, we have (at least) $3$ saddles, $3$ maxima and $3$ minima which gives at least $10$ critical points since the number of critical points is even.
4. Proof of Proposition 2
Since the proof is quite technical, we recommend the reader to start with proving the proposition for skew triangles only, using the following:
Lemma 9.
[5]
A spherical triangle $ABC$ is skew iff (up to relabeling of the vertices), one has $|CA|>\pi/2$, $|BC|<\pi/2$, $\angle ABC>\pi/2$, $\angle BAC<\pi/2$, $\angle BCA<\pi/2$, $|BA|>\pi/2$, $|AZ|>\pi/2$ for the point $Z\in AC$ such that $ZB$ is orthogonal to $BC$.
Throughout the section $P$ is a generic skew spherical polygon.
Definition 6.
A bigon is a spherical polygon with two edges. It is bounded by two halves of big circles. The minimal bigon $Big(P)$ of a spherical polygon $P$ is the bigon containing $P$ which is minimal by inclusion, or, equivalently, with the minimal angle between its edges.
Lemma 10.
$P$ intersects each edge of its minimal bigon $Big(P)$ by a segment whose interior contains the midpoint of the edge, see Fig 4.
Proof.
Denote the edges of $Big(P)$ by $l_{1}$ and $l_{2}$, and their midpoints by $R$ and $T$. Firstly, prove that $R$ and $T$ lie on $\partial P$. Indeed, if $R\notin\partial P$ a rotation of $l_{1}$ decreases the bigon.
Figure 4. $P$ intersects $Big(P)$ by two segments, containing the midpoints $R$ and $T$.
Next, prove that $P$ intersects each edge of $Big(P)$ by a segment. If not, two cases are possible:
(1) $P$ intersects each of the edges by a point. Then a rotation decreases $Big(P)$.
(2) $P$ intersects one edge by a point, and the other edge by a segment. Then necessarily the intersection point is the midpoint of the edge, say, $R$, and also the intersection segment contains $T$. Let us show that $P$ is not skew. Consider the segment $RT$; it lies in $P$ and cuts the bigon into two equal halves. Let $X$ be the midpoint of $RT$, then $X\in Core(P)$ and $SQD^{sph}_{X}$ has a short maximum at $R$ and a short minimum at $T$. Since $|XR|=|XT|$, $SQD^{sph}_{X}$ has additionally at least two short maxima lying in different halves of bigon. Altogether we have 3 short maxima. A contradiction.
Corollary 1.
The midpoint $X$ of $Big(P)$ lies in $P$. Besides, $X\in Core(P)$ and $SQD^{sph}_{X}$ has two (short) minima at $R$ and $T$.
Proposition 3.
If $P$ is skew, then $Big(P)$ is acute-angled.
Proof.
Assume the contrary: $Big(P)$ is obtuse-angled. Then all the angles of $P$ are also obtuse (this follows from minimality).
The proof is based on two lemmata.
A spherical segment $AB$ is short if $|AB|<\pi/2$. Otherwise, the segment is called long.
Definition 7.
Let $A$ be a vertex of $P$. Set $\mathcal{AR}^{spher}(A)$ be the set of points $Y\subset P$ such that $A$ is a short local maximum of $SQD_{Y}^{spher}$.
Lemma 11.
If $P$ is skew and $Big(P)$ is obtuse-angled, then no two short edges of $P$ share a vertex.
Proof.
Let short edges $AB$ and $AC$ be incident. Let the point $Z$ be as is depicted in Fig. 5. That is, $Z$ is the intersection of two perpendiculars to the edges emanating from $B$ and $C$. We have: $|ZB|<\pi/2,|ZA|<\pi/2,|ZC|<\pi/2,Z\in\mathcal{AR}^{spher}(A)$. If $Z\in P$, then $P$ is not skew, since for point $Y$ lying close to $Z$, $SQD^{sph}_{Y}$ has $3$ short maxima. Indeed, shift $Z$ a little bit upwards.
Figure 5. The point $Z$ and the line $l$ for the proof of Lemma 11
Assume that $Z\notin P$. Then there is a (geodesic) line $l$ that separates $Z$ and $P$. Consider the intersection points $D$ and $E$ of $l$ with the sides of the angle $A$.
(1) If $D$ and $E$ lie in the upper hemisphere, then $P$ is contained in a quarter of the sphere, so the minimal bigon cannot be obtuse-angled.
(2) Let $D$ lie in the bottom hemisphere. All angles in the triangle $ADE$ are obtuse, therefore the intersection point of $l$ with the polar line of $A$ doesn’t lie in $\mathcal{AR}^{spher}(A)$. Then $Z\notin\mathcal{AR}^{spher}(A)$. A contradiction. ∎
Lemma 12.
If $P$ is skew and $Big(P)$ is obtuse-angled, then $P$ has at most $4$ vertices.
Proof.
The midline $RT$ cuts the bigon into two halves. If $P$ has at least 5 vertices, then by Lemma 11 there exists a long edge $AB$, contained in one half of the bigon. For the midpoint $X$ of the bigon, we have: $|XA|<\pi/2,|XB|<\pi/2,|AB|>\pi/2$. Then $\angle XAB$ and $\angle XBA$ are acute, so $X$ projects on $AB$. Then $SQD^{sph}_{X}$ has at least 3 minima. Since $X\in Core(P)$, the function $SQD^{sph}_{X}$ has at least 3 short maxima, so $P$ is not skew.
It remains to analyze triangled and quadrilaterals.
  1. (1)
    $P$ is a triangle. All its angles are obtuse, so it’s not skew by Lemma 9.
  2. (2)
    $P$ is a quadrilateral. Consider the triangle $KLM$ as in Fig.6. It is a connected component of $Big(P)\setminus P$. We have: $|KL|<\pi/2$, $|KM|<\pi/2$, $|LM|<\pi/2$, $\angle KLM<\pi/2$, $\angle KML<\pi/2$, $\angle LKM>\pi/2$. Then spherical geometry tells us that $K$ projects to $LM$. Let $KH\perp LM$. Moreover, $\angle LKH<\pi/2$, $\angle MKH<\pi/2$. Let $S=KH\cap RT$. Then $|SR|<\pi/2$, $|ST|<\pi/2$. Therefore $S\in Core(P)$, and $SQD^{sph}_{S}$ has at least three (short) minima at $R,T,$ and $H$. Therefore there are three short maxima, so $P$ is not skew.
Proposition 3 is proven. ∎
Figure 6. Case 2: $P$ is a quadrilateral.
Proposition 3 implies that the midline $RT$ is contained in $Core(P)$. With this knowledge we proceed with the proof of Proposition 2. So let $X$ be the midpoint of the minimal bigon $Big(P)$.
Denote by $M_{1}$ and $M_{2}$ maxima points of $SQD^{sph}_{X}$. They lie in different halves of the bigon.
Lemma 13.
$M_{1}$ and $M_{2}$ are acute vertices.
Proof.
Assume that $M_{1}$ obtuse. Two cases are possible:
  1. (1)
    $M_{1}$ lies on an edge of the bigon. $M_{1}$ is clearly not a vertex of bigon since the bigon is acute-angled. Let a point $X$ travels along the midline in the direction of the edge containing $M_{1}$. The points $R$ and $T$ are short minima of $SQD^{sph}_{X}$ all the time. Since these minima are fixed, and we have not other minima, there are no bifurcations on the way. But if $X$ is close to $R$, and $M_{1}$ is an obtuse vertex, it can not be maximum of $SQD^{sph}_{X}$.
  2. (2)
    $M_{1}$ lies in the interior of the bigon. Then incident edges $M_{1}A$ and $M_{1}B$ are short, and angle between them is obtuse. Consider the point $Z$ which is defined as the intersection of two perpendiculars, as in Fig. 7. Clearly, $|ZB|<\pi/2,|ZM_{1}|<\pi/2,|ZA|<\pi/2,Z\in\mathcal{AR}^{spher}(M_{1})$. Also $X\in\mathcal{AR}^{spher}(M_{1})$. We also have $Z\in conv(A,B,M_{1},X)$. Then $Z\in P$, and $P$ is not skew.
Figure 7. For the proof of Lemma 13
Clearly, there are no more acute vertices. Indeed, if a vertex is acute, it’s a short maximal point of $SQD^{sph}_{X}$. Proposition 2, (1) is proven.
Lemma 14.
Let $A$ be a vertex of $P$, $p\in P$ and $|M_{1}p|<\pi/2$, $|M_{2}p|<\pi/2$ ($M_{1}$ and $M_{2}$ are the acute vertices). Then $|Ap|<\pi/2$.
Proof.
Assume that $A$ lies in the same half of bigon as $M_{1}$. Then we have:$|M_{1}p|<\pi/2$, $|M_{1}A|\leqslant\pi/2$, $\angle AM_{1}p<\pi/2$, which implies $|Ap|<\pi/2$. ∎
The following lemma completes the proof of Proposition 2.
Lemma 15.
$|M_{1}M_{2}|>\pi/2$.
Proof.
Assume the contrary, that is, $|M_{1}M_{2}|<\pi/2$. Apply twice Lemma 14 for $p=M_{1}$ and $p=M_{2}$ one by one. Then for any vertex $A$ we have: $|M_{1}A|<\pi/2,|M_{2}A|<\pi/2$. Select a point $A^{\prime}$ near a vertex $A$, inside its spherical active region. So, $SQD^{sph}_{A^{\prime}}$ has three short maxima at $M_{1},M_{2},$ and $A$. Therefore $P$ is not skew. A contradiction. ∎